# Whitehead’s problem and condensed mathematics

### v2

@@ 5. Uncountable groups
In what follows, given a cardinal $\kappa$, $2^{\kappa}$ denotes the topological space consisting of all functions $f:\kappa\rightarrow 2$, equipped with the product topology. In particular, basic clopen subsets of $2^{\kappa}$ are of the form $U_{\sigma}:=\{f\in 2^{\kappa}\mid\sigma\subseteq f\}$, where $\sigma$ is a finite partial function from $\kappa$ to $2$. Note that $2^{\kappa}$ is a profinite set.
@@
Thus, for the rest of the proof we may assume that $A$ is almost free and $|F|=\kappa$. The case of $\kappa=\aleph_{0}$ is covered by Theorem 4.2, so we may assume that $\kappa>\aleph_{0}$. Let $\vec{M}=\langle M_{\alpha}\mid\alpha<\kappa\rangle$ be a $\in$-increasing, continuous sequence of elementary submodels of $(H(\chi),\in,\vartriangleleft,A,F,K)$ such that $|M_{\alpha}|<\kappa$ and $\delta_{\alpha}:=M_{\alpha}\cap\kappa\in\kappa$ for all $\alpha<\kappa$. For each $\alpha<\kappa$, let $A_{\alpha}:=A\cap M_{\alpha}$, $F_{\alpha}:=F\cap M_{\alpha}$, and $K_{\alpha}:=K\cap M_{\alpha}$. Since $A$, $F$, and $K$ each have size at most $\kappa$ and $\vec{M}$ is $\in$-increasing, we have $A=\bigcup_{\alpha<\kappa}A_{\alpha}$, $F=\bigcup_{\alpha<\kappa}F_{\alpha}$, and $K=\bigcup_{\alpha<\kappa}K_{\alpha}$.
Let $E$ be the set of $\alpha<\kappa$ such that there exists $\beta$ such that $\alpha<\beta<\kappa$ and $A_{\beta}/A_{\alpha}$ is nonfree. Note that, if $\alpha<\beta<\beta^{\prime}<\kappa$ and $A_{\beta}/A_{\alpha}$ is nonfree, then also $A_{\beta^{\prime}}/A_{\alpha}$ is nonfree, since it contains $A_{\beta}/A_{\alpha}$ as a subgroup. We claim that $E$ is stationary. If not, then there exists a club $D\subseteq\kappa$ such that, for all $\alpha<\beta$, both in $D$, the group $A_{\beta}/A_{\alpha}$ is free. Moreover, since $A$ is almost free, we know that $A_{\min(D)}$ is free. Therefore, it is straightforward to build a basis for $A$ by recursion along $D$ (cf. [5, Theorem 5.3]), contradicting the assumption that $A$ is not free.
@@
For each $\alpha\in S$, let $A_{\alpha}^{*}:=A_{\alpha+1}/A_{\alpha}$, $F_{\alpha}^{*}:=F_{\alpha+1}/F_{\alpha}$, and $K_{\alpha}^{*}:=K_{\alpha+1}/K_{\alpha}$. Similarly, let $B_{F,\alpha}^{*}:=B_{F}\cap(M_{\alpha+1}\setminus M_{\alpha})$ and $B_{K,\alpha}^{*}:=B_{K}\cap(M_{\alpha+1}\setminus M_{\alpha})$. Note that each of $F_{\alpha}^{*}$ and $K_{\alpha}^{*}$ is free, with bases given by $\{z+F_{\alpha}\mid z\in B_{F,\alpha}^{*}\}$ and $\{t+K_{\alpha}\mid t\in B_{K,\alpha}^{*}\}$, respectively. Also, our free resolution of $A$ induces a free resolution
@@ 6. Cohen forcing
Before proving Theorem 6.1, we make the following observation.
@@
Note that this construction continues to make sense in any outer model of $V_{1}$. Let $\dot{g}$ be the canonical $\mathrm{Add}(\omega,\kappa)$-name for the union of the generic filter, and note that $\dot{g}$ is a name for a function from $\kappa$ to $\{0,1\}$. Then let $\dot{\varphi}$ be an $\mathrm{Add}(\omega,\kappa)$-name for $\varphi(\dot{g})$. We claim that $\dot{\varphi}$ is forced to be as desired. To see this, let $\dot{\mathbb{P}}$ be an $\mathrm{Add}(\omega,\kappa)$-name for a ccc forcing notion, and suppose for sake of contradiction that $\dot{\psi}$ is an $\mathrm{Add}(\omega,\kappa)\ast\dot{\mathbb{P}}$-name that is forced to be an element of $\mathrm{Hom}(F,\mathbb{Z})$ extending $\dot{\varphi}$.
@@ 7. Stone spaces
Given an infinite cardinal $\kappa$, let $\mathbb{B}_{\kappa}$ denote the Boolean completion of $\mathrm{Add}(\omega,\kappa)$, and let $S_{\kappa}$ denote the Stone space $S(\mathbb{B}_{\kappa})$. We will sometimes slightly abuse notation and think of $\mathrm{Add}(\omega,\kappa)$ as a subset of $\mathbb{B}_{\kappa}$. Given sets $X$ and $Y$, we now highlight a translation between $\mathbb{B}_{\kappa}$-names for functions from $X$ to $Y$ and continuous functions from $S_{\kappa}$ to the product space $Y^{X}$. Recall that a clopen basis for $S_{\kappa}$ is given by all sets of the form
@@
    Suppose now that $Y$ is a finite set and $\dot{f}$ is a $\mathbb{B}_{\kappa}$-name for a function from $X$ to $Y$. Define a function $\varphi_{\dot{f}}:S_{\kappa}\rightarrow Y^{X}$ as follows: for each $U\in S_{\kappa}$ and $x\in X$, let $\varphi_{\dot{f}}(x)$ be the unique $y\in Y$ such that $\llbracket\dot{f}(x)=y\rrbracket_{\mathbb{B}_{\kappa}}\in U$. Note that such a $y$ must exist, since $\{\llbracket\dot{f}(x)=y\rrbracket_{\mathbb{B}_{\kappa}}\mid y\in Y\}$ is a finite maximal antichain in $\mathbb{B}_{\kappa}$ and therefore must intersect $U$.
@@
We claim that $\varphi$ is as desired. We clearly have $\varphi(U)[B_{K}]\subseteq\{0,1\}$ for all $U\in S_{\kappa}$. Suppose for sake of contradiction that $\psi:S_{\kappa}\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ is a continuous map and extends $\varphi$ pointwise. Let $\dot{f}_{\psi}$ be the $\mathbb{B}_{\kappa}$-name for a function from $F$ to $\mathbb{Z}$ given by Definition 7.1(1).
@@
Suppose for sake of contradiction that there are $y,z\in F$ and $b\in\mathbb{B}_{\kappa}^{+}$ such that $b\Vdash_{\mathbb{B}_{\kappa}}\dot{f}_{\psi}(y)+\dot{f}_{\psi}(z)\neq\dot{f}_{\psi}(y+z)$. By extending $b$, we may assume that there are $k_{y}$, $k_{z}$, and $k_{yz}$ such that $b\Vdash_{\mathbb{B}_{\kappa}}(\dot{f}_{\psi}(y),\dot{f}_{\psi}(z),\dot{f}_{\psi}(y+z))=(k_{y},k_{z},k_{yz})$ and $k_{y}+k_{z}\neq k_{yz}$. Let $U\in S_{\kappa}$ be such that $b\in U$. Then by the definition of $\dot{f}_{\psi}$, we have $\psi(U)(y)=k_{y}$, $\psi(U)(z)=k_{z}$, and $\psi(U)(y+z)=k_{yz}$, contradicting the fact that $\psi(U)\in\mathrm{Hom}(F,\mathbb{Z})$. ∎
@@
Suppose for sake of contradiction that there is $t\in B_{K}$ and $b\in\mathbb{B}_{\kappa}^{+}$ such that $b\Vdash_{\mathbb{B}_{\kappa}}\dot{f}_{\psi}(t)\neq\dot{\tau}(t)$. By extending $b$ if necessary, we may assume without loss of generality that there are $k_{\psi},k_{\tau}\in\mathbb{Z}$ such that $b\Vdash_{\mathbb{B}_{\kappa}}(\dot{f}_{\psi}(t),\dot{\tau}(t))=(k_{\psi},k_{\tau})$. Let $U\in S_{\kappa}$ be such that $b\in U$. Then by the definition of $\varphi_{\dot{f}}$, we have $\varphi_{\dot{f}}(U)(t)=\varphi(U)(t)=k_{\psi}$, and by the definition of $\dot{f}_{\psi}$, we have $\psi(U)(t)=k_{\tau}$, contradicting the fact that $\psi$ extends $\varphi$ pointwise. ∎
@@ 8. Topological spaces in forcing extensions
We end this section with a few remarks. First, as was mentioned above, if $X$ is a locally compact Hausdorff space, then $\hat{X}_{\mathcal{K}}$ is the canonical interpretation of $X$ in $V^{\mathbb{B}}$ in the sense of [29]. Recall also that the category $\mathsf{ED}$ is precisely the category of Stone spaces of complete Boolean algebras; recall as well that $\underline{X}(S)$ equals, by definition, $C(S,X)$ for any $S\in\mathsf{ED}$. Therefore, the results of this section show that one way to think of the embedding of the category of locally compact Hausdorff spaces into $\mathsf{Cond}$ is to recognize that, for all such spaces $\underline{X}$, the condensed set $\underline{X}$ is precisely an organized presentation of all forcing names for points in the canonical interpretations of $X$ in all possible set forcing extensions.
Second, we note that the operation $X\mapsto\hat{X}_{\mathcal{K}}$ commutes with products. Namely, if $X_{i}$ is a Hausdorff space for all $i\in I$, then $(\widehat{\prod_{i\in I}X_{i}})_{\mathcal{K}}\cong\prod_{i\in I}(\hat{X}_{i})_{\mathcal{K}}$. This is because the collection
@@
Third, even though Theorem 8.11 holds for all Hausdorff spaces $X$, there is an important sense in which the correspondence is meaningfully stronger in case $X$ is compact. In particular, let $\dot{k}$ be the $\mathbb{B}$-name for the homeomorphism we constructed to witness Theorem 8.11. If $X$ is compact, then, for every $\mathbb{B}$-name $\dot{x}$ for an element of $\hat{X}_{\mathcal{K}}(=\hat{X})$, there is a $g\in C(S,X)$ such that $1_{\mathbb{B}}\Vdash\dot{k}([g])=\dot{x}$. In particular, in this case we can take $\dot{x}$ to be a $\mathbb{B}$-name for a maximal filter on $\tau_{c}$, then then let $g=f_{\dot{x}}$, as defined at the outset of this section. However, in general this may not be the case, and there may be $\mathbb{B}$-names $\dot{x}$ for elements of $\hat{X}_{\mathcal{K}}$ for which there is no $g\in C(S,X)$ such that $1_{\mathbb{B}}\Vdash\dot{k}([g])=\dot{x}$. For such $\dot{x}$, one may need to work below some condition $b\in\mathbb{B}$ that forces $\dot{x}$ to lie inside the interpretation of some particular $K\in\mathcal{K}(X)$; once one does so, one may then find $g\in C(S,X)$ such that $b\Vdash\dot{k}([g])=\dot{x}$. This issue will lead to some slight complications and asymmetries in the next section.
@@ 9. Compact interpretations and continuous functions
Now suppose that $V=L$, and let $A$ be an abelian group of size $\aleph_{1}$ that satisfies Chase’s condition but is not free. For concreteness, suppose it is constructed as in [5, Theorem 7.3]. It is readily verified that this construction is absolute between $V$ and outer models with the same $\aleph_{1}$; in particular, in any ccc forcing extension, it remains true that $A$ satisfies Chase’s condition and is not free. Let $\mathbb{B}$ be the Boolean completion of the standard ccc forcing iteration to force $\sf MA_{\omega_{1}}$, and let $0\rightarrow K\rightarrow F\rightarrow A\rightarrow 0$ be a free resolution of $A$. Since $V=L$, $A$ is not Whitehead in $V$, so there is a map $\tau\in\mathrm{Hom}(K,\mathbb{Z})$ that does not lift to a map $\sigma\in\mathrm{Hom}(F,\mathbb{Z})$. Then the constant map $\varphi:S(\mathbb{B})\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ taking value $\tau$ witnesses that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S(\mathbb{B}))\neq 0$. However, since $A$ still satisfies Chase’s condition in $V^{\mathbb{B}}$ and $\sf MA_{\omega_{1}}$ holds there, it follows that $\mathbb{B}$ forces that $A$ is Whitehead.

### v3

@@ 5. Uncountable groups
In what follows, given a cardinal $\kappa$, $2^{\kappa}$ denotes the topological space consisting of all functions $f:\kappa\rightarrow 2$, equipped with the product topology. In particular, basic clopen subsets of $2^{\kappa}$ are of the form $U_{\sigma}:=\{f\in 2^{\kappa}\mid\sigma\subseteq f\}$, where $\sigma$ is a finite partial function from $\kappa$ to $2$. Note that $2^{\kappa}$ is a profinite set. Our next result establishes item (1) of Theorem 1.3.
@@
Thus, for the rest of the proof we will assume that $A$ is almost free and $|F|=\kappa$. The case of $\kappa=\aleph_{0}$ is covered by Theorem 4.2, so we may assume that $\kappa>\aleph_{0}$. Let $\vec{M}=\langle M_{\alpha}\mid\alpha<\kappa\rangle$ be a $\in$-increasing, continuous sequence of elementary submodels of $(H(\chi),\in,\vartriangleleft,A,F,K)$ such that $|M_{\alpha}|<\kappa$ and $\delta_{\alpha}:=M_{\alpha}\cap\kappa\in\kappa$ for all $\alpha<\kappa$. For each $\alpha<\kappa$, let $A_{\alpha}:=A\cap M_{\alpha}$, $F_{\alpha}:=F\cap M_{\alpha}$, and $K_{\alpha}:=K\cap M_{\alpha}$. Since $A$, $F$, and $K$ each have size at most $\kappa$ and $\vec{M}$ is $\in$-increasing, we have $A=\bigcup_{\alpha<\kappa}A_{\alpha}$, $F=\bigcup_{\alpha<\kappa}F_{\alpha}$, and $K=\bigcup_{\alpha<\kappa}K_{\alpha}$.
Let $E$ be the set of $\alpha<\kappa$ such that there exists a $\beta$ such that $\alpha<\beta<\kappa$ and $A_{\beta}/A_{\alpha}$ is nonfree. Note that, if $\alpha<\beta<\beta^{\prime}<\kappa$ and $A_{\beta}/A_{\alpha}$ is nonfree, then also $A_{\beta^{\prime}}/A_{\alpha}$ is nonfree, since it contains $A_{\beta}/A_{\alpha}$ as a subgroup. We claim that $E$ is stationary. If not, then there exists a club $D\subseteq\kappa$ such that, for all $\alpha<\beta$, both in $D$, the group $A_{\beta}/A_{\alpha}$ is free. Moreover, since $A$ is almost free, we know that $A_{\min(D)}$ is free. Therefore, it is straightforward to build a basis for $A$ by recursion along $D$ (cf. [9, Theorem 5.3]), contradicting the assumption that $A$ is not free.
@@
For each $\alpha\in S$, let $A_{\alpha}^{*}:=A_{\alpha+1}/A_{\alpha}$, $F_{\alpha}^{*}:=F_{\alpha+1}/F_{\alpha}$, and $K_{\alpha}^{*}:=K_{\alpha+1}/K_{\alpha}$. Similarly, let $B_{F,\alpha}^{*}:=B_{F}\cap(M_{\alpha+1}\setminus M_{\alpha})$ and $B_{K,\alpha}^{*}:=B_{K}\cap(M_{\alpha+1}\setminus M_{\alpha})$. Note that each $F_{\alpha}^{*}$ and $K_{\alpha}^{*}$ is free, with bases given by $\{z+F_{\alpha}\mid z\in B_{F,\alpha}^{*}\}$ and $\{t+K_{\alpha}\mid t\in B_{K,\alpha}^{*}\}$, respectively. Also, our free resolution of $A$ induces a free resolution
@@ 6. Cohen forcing
Note that this establishes item (2) of Theorem 1.3. Before proving Theorem 6.1, we make the following observation.
@@
Note that this construction continues to make sense in any outer model of $V_{1}$. Let $\dot{g}$ be the canonical $\mathrm{Add}(\omega,\kappa)$-name for the union of the generic filter, and note that $\dot{g}$ is a name for a function from $\kappa$ to $\{0,1\}$. Then let $\dot{\varphi}$ be an $\mathrm{Add}(\omega,\kappa)$-name for $\varphi(\dot{g})$. We claim that $\dot{\varphi}$ is forced to be as desired. To see this, let $\dot{\mathbb{P}}$ be an $\mathrm{Add}(\omega,\kappa)$-name for a ccc forcing notion, and suppose for the sake of contradiction that $\dot{\psi}$ is an $\mathrm{Add}(\omega,\kappa)\ast\dot{\mathbb{P}}$-name that is forced to be an element of $\mathrm{Hom}(F,\mathbb{Z})$ extending $\dot{\varphi}$.
@@ 7. Stone spaces
Given an infinite cardinal $\kappa$, let $\mathbb{B}_{\kappa}$ denote the Boolean completion of $\mathrm{Add}(\omega,\kappa)$, and let $S_{\kappa}$ denote the Stone space of $\mathbb{B}_{\kappa}$. We will sometimes slightly abuse notation and think of $\mathrm{Add}(\omega,\kappa)$ as a subset of $\mathbb{B}_{\kappa}$. Given sets $X$ and $Y$, we now highlight a translation between $\mathbb{B}_{\kappa}$-names for functions from $X$ to $Y$ and continuous functions from $S_{\kappa}$ to the product space $Y^{X}$. Recall that a clopen basis for $S_{\kappa}$ is given by all sets of the form
@@
    Suppose now that $Y$ is a finite set and $\dot{f}$ is a $\mathbb{B}_{\kappa}$-name for a function from $X$ to $Y$. Define a function $\varphi_{\dot{f}}:S_{\kappa}\rightarrow Y^{X}$ as follows: for each $U\in S_{\kappa}$ and $x\in X$, let $\varphi_{\dot{f}}(U)(x)$ be the unique $y\in Y$ such that $\llbracket\dot{f}(x)=y\rrbracket_{\mathbb{B}_{\kappa}}\in U$. Note that such a $y$ must exist, since $\{\llbracket\dot{f}(x)=y\rrbracket_{\mathbb{B}_{\kappa}}\mid y\in Y\}$ is a finite maximal antichain in $\mathbb{B}_{\kappa}$ and therefore must intersect $U$.
@@
We claim that $\varphi$ is as desired. We clearly have $\varphi(U)[B_{K}]\subseteq\{0,1\}$ for all $U\in S_{\kappa}$. Suppose for the sake of contradiction that $\psi:S_{\kappa}\rightarrow\mathrm{Hom}(F,\mathbb{Z})$ is a continuous map and extends $\varphi$ pointwise. Let $\dot{f}_{\psi}$ be the $\mathbb{B}_{\kappa}$-name for a function from $F$ to $\mathbb{Z}$ given by Definition 7.1(1).
@@
Suppose for the sake of contradiction that there are $y,z\in F$ and $b\in\mathbb{B}_{\kappa}^{+}$ such that $b\Vdash_{\mathbb{B}_{\kappa}}\dot{f}_{\psi}(y)+\dot{f}_{\psi}(z)\neq\dot{f}_{\psi}(y+z)$. By extending $b$, we may assume that there are $k_{y}$, $k_{z}$, and $k_{yz}$ such that $b\Vdash_{\mathbb{B}_{\kappa}}(\dot{f}_{\psi}(y),\dot{f}_{\psi}(z),\dot{f}_{\psi}(y+z))=(k_{y},k_{z},k_{yz})$ and $k_{y}+k_{z}\neq k_{yz}$. Let $U\in S_{\kappa}$ be such that $b\in U$. Then by the definition of $\dot{f}_{\psi}$, we have $\psi(U)(y)=k_{y}$, $\psi(U)(z)=k_{z}$, and $\psi(U)(y+z)=k_{yz}$, contradicting the fact that $\psi(U)\in\mathrm{Hom}(F,\mathbb{Z})$. ∎
@@
Suppose for the sake of contradiction that there is $t\in B_{K}$ and $b\in\mathbb{B}_{\kappa}^{+}$ such that $b\Vdash_{\mathbb{B}_{\kappa}}\dot{f}_{\psi}(t)\neq\dot{\tau}(t)$. By extending $b$ if necessary, we may assume without loss of generality that there are $k_{\psi},k_{\tau}\in\mathbb{Z}$ such that $b\Vdash_{\mathbb{B}_{\kappa}}(\dot{f}_{\psi}(t),\dot{\tau}(t))=(k_{\psi},k_{\tau})$. Let $U\in S_{\kappa}$ be such that $b\in U$. Then by the definition of $\varphi_{\dot{f}}$, we have $\varphi_{\dot{f}}(U)(t)=\varphi(U)(t)=k_{\psi}$, and by the definition of $\dot{f}_{\psi}$, we have $\psi(U)(t)=k_{\tau}$, contradicting the fact that $\psi$ extends $\varphi$ pointwise. ∎
@@ 8. Topological spaces in forcing extensions
We end this section with a few remarks. First, as was mentioned above, if $X$ is a locally compact Hausdorff space, then $\hat{X}_{\mathcal{K}}$ is the canonical interpretation of $X$ in $V^{\mathbb{B}}$ in the sense of [36]. Recall also that the category $\mathsf{ED}$ is precisely the category of Stone spaces of complete Boolean algebras, and that $\underline{X}(S)$ equals, by definition, $C(S,X)$ for any $S\in\mathsf{ED}$. Therefore, the results of this section show that one way to think about the embedding of the category of locally compact Hausdorff spaces into $\mathsf{Cond}$ is to recognize that, as noted, for all such spaces $X$, the condensed set $\underline{X}$ is simply an organized presentation of all forcing names for points in the canonical interpretations of $X$ in all possible set forcing extensions.
Second, write $cg$ for the left adjoint of the inclusion of the category of compactly generated topological spaces into the category of topological spaces; recall also that the embedding $X\mapsto\underline{X}$ of the category of $T_{1}$ topological spaces into Cond possesses a left adjoint whose counit is given by $X\mapsto cg(X)$ (modulo a small technical point; see [26, pp. 8-9 and Appendix to Lecture II]). Observe next that in the $\mathbb{B}=\{0,1\}$ (and hence $S(\mathbb{B})=*$) instance of Theorem 8.11, $\hat{X}_{\mathcal{K}}$ is homeomorphic to $X$. Had we endowed it instead with the $\hat{\tau}^{\lim}_{\mathcal{K}}$ topology considered above, it would be homeomorphic to $cg(X)$, and the $\hat{\tau}^{\lim}_{\mathcal{K}}$ topology corresponds more generally to that given by post-composing the conversion recorded in Theorem 8.11 with the functor $cg$.
Third, we note that the operation $X\mapsto\hat{X}_{\mathcal{K}}$ commutes with products. Namely, if $X_{i}$ is a Hausdorff space for every $i\in I$, then $(\widehat{\prod_{i\in I}X_{i}})_{\mathcal{K}}\cong\prod_{i\in I}(\hat{X}_{i})_{\mathcal{K}}$. This is because the collection
@@
Fourth, even though Theorem 8.11 holds for all Hausdorff spaces $X$, there is an important sense in which the correspondence is meaningfully stronger in case $X$ is compact. In particular, let $\dot{k}$ be the $\mathbb{B}$-name for the homeomorphism we constructed to witness Theorem 8.11. If $X$ is compact, then, for every $\mathbb{B}$-name $\dot{x}$ for an element of $\hat{X}_{\mathcal{K}}(=\hat{X})$, there is a $g\in C(S,X)$ such that $1_{\mathbb{B}}\Vdash\dot{k}([g])=\dot{x}$. In particular, in this case we can take $\dot{x}$ to be a $\mathbb{B}$-name for a maximal filter on $\tau_{c}$, then then let $g=f_{\dot{x}}$, as defined at the outset of this section. However, in general this may not be the case, and there may be $\mathbb{B}$-names $\dot{x}$ for elements of $\hat{X}_{\mathcal{K}}$ for which there is no $g\in C(S,X)$ such that $1_{\mathbb{B}}\Vdash\dot{k}([g])=\dot{x}$. For such $\dot{x}$, one may need to work below some condition $b\in\mathbb{B}$ that forces $\dot{x}$ to lie inside the interpretation of some particular $K\in\mathcal{K}(X)$; once one does so, one may then find $g\in C(S,X)$ such that $b\Vdash\dot{k}([g])=\dot{x}$. This issue will lead to some slight complications and asymmetries in the next section.
@@ 9. Compact interpretations and continuous functions
Now suppose that $V=L$, and let $A$ be an abelian group of size $\aleph_{1}$ that satisfies Chase’s condition but is not free. For concreteness, suppose it is constructed as in [9, Theorem 7.3]. It is readily verified that this construction is absolute between $V$ and outer models with the same $\aleph_{1}$; in particular, in any ccc forcing extension, it remains true that $A$ satisfies Chase’s condition and is not free. Let $\mathbb{B}$ be the Boolean completion of the standard ccc forcing iteration to force $\sf MA_{\omega_{1}}$, and let $0\rightarrow K\rightarrow F\rightarrow A\rightarrow 0$ be a free resolution of $A$. Since $V=L$, $A$ is not Whitehead in $V$, so there is a map $\tau\in\mathrm{Hom}(K,\mathbb{Z})$ that does not lift to a map $\sigma\in\mathrm{Hom}(F,\mathbb{Z})$. Then the constant map $\varphi:S(\mathbb{B})\rightarrow\mathrm{Hom}(K,\mathbb{Z})$ taking value $\tau$ witnesses that $\underline{\mathrm{Ext}}^{1}_{\mathsf{CondAb}}(\underline{A},\underline{\mathbb{Z}})(S(\mathbb{B}))\neq 0$. However, since $A$ still satisfies Chase’s condition in $V^{\mathbb{B}}$ and $\sf MA_{\omega_{1}}$ holds there, it follows that $\mathbb{B}$ forces that $A$ is Whitehead.